How do you convert something with a radius of 10 feet that makes 6 revolutions/minute to inches per minute? Thank you
Answer by Edwin McCravy(20056)

1 revolution's distance in feet = 2*pi*r = 2*pi*10 ft = 20pi ft.

1 revolution's distance in inches = 20pi*12 = 240pi inches

6 revolution's distance in inches = 1440pi inches

6 revolution's distance in inches/minute = 1440pi inches/minute

Edwin